1. Write an equation of the line containing (5,2) and (7,-4). Express the answer in standard form with integer coefficients.
2. Write an equation of the line containing (4,-1) and (-2,-6). Express the answer in standard form with integer coefficients.

Answer by Boreal(15235)

slope is -6/2=-3
point slope formula y-y1=m(s-x1), m slope and (x1, y1) point
y-2=-3(x-5)
y=-3x+17
3x+y=17
the second one
slope is 5/6
y+1=(5/6)(x-4)=(5/6)x-20/6
y=(5/6)x-26/6
multiply by 6 and rearrange
6y-5x=-26
5x-6y=26